An airplane flew for 4 hours with a 20-km/h tail wind. The return flight against the same wind took 5 hours. Find the speed of the plane still in air.
Answer by mananth(16949)

Plane speed x km/h
wind speed 20 km/h
..
against currentspeed x- 20
withcurrent x+ 20
..
Against current 5 hours
with current x+y 4 hours
Distance is same. D=r*t
5 ( x - 20 ) = 4 ( x + 20 )
5 x - 100 = 4 x + 80
5 x - 4 x = 100 + 80
1 x = 180
/ 1
x = 180 km/h
Plane speed = 180 km/h